1. A compound has 45 g of iron, 6 g of oxygen, and 3 g of sulfur. Find the percent of each element.


2. A compound has 3 g of sodium and 5 g of bromide. Find the percent of each.


1
Expert's answer
2021-05-11T05:56:48-0400

1) Formula of percentage composition willl be -


Mass % ="\dfrac{mass}{total\ mass}\times 100" %


Mass % of iron = "\dfrac{45}{54}\times 100=83.33"


Mass % of oxygen ="\dfrac{6}{54}\times 100=11.11"


Mass % of sulphur = "\dfrac{3}{54}\times 100=5.55"


2) Mass % ="\dfrac{mass}{total\ mass}\times 100" %


Mass % of sodium ="\dfrac{3}{8}\times 100=37.5"


Mass % of bromide = "\dfrac{5}{8}\times 100=62.5"